# Noncommutative metric geometry of quantum circle bundles

### v1

@@ 1. Introduction
In the present paper we construct what could be called a “twisted spectral quadruple” for each of the higher quantum spheres. The key point is here that most of the quantized enveloping algebra is kept in its deformed state except for the element $K_{r}$ which is, in some sense, replaced by its classical counterpart. The element $K_{r}$ plays a special role since it is related to the circle action which describes the quantum circle bundle structure of the quantum sphere, singling out quantum projective space as invariant elements. Replacing $K_{r}$ with its classical counterpart ensures that the compact resolvent condition is satisfied for our spectral data but the price is relatively high since we are sometimes forced to consider a pair of abstract Dirac operators instead of a single one. This is precisely because of the commutator condition: One of our abstract Dirac operators interacts with the coordinate algebra via straight bounded commutators whereas the other one interacts via twisted bounded commutators. Nonetheless, both of these Dirac operators are needed in order to describe the spectral geometry of the quantum sphere since only their sum has compact resolvent, not the individual factors. Our approach is therefore related to the papers [KRS12, KaKy22, Ć24] and our work do indeed generalize parts of the results in [KaKy22].
@@
As already mentioned, we are not only focusing on the quantum spheres since our main results are formulated in the generality of quantum circle bundles. In particular, we provide general conditions which allow us to write down “twisted spectral quadruples” for quantum circle bundles. A part from the quantum spheres our methods might therefore also apply to the “quantum Grassmann spheres” which are quantum circle bundles over the quantum Grassmannians, [TaTo91, Fio99, CMOB23]. Treating this case in detail would however lead us to far astray for the time being and we therefore leave it open for future investigations. It should also be emphasized that our results to some extend cover the results of [BMR10, HSWZ13, KaKy21] regarding crossed products by the integers as one may indeed view the corresponding algebraic crossed product as a quantum circle bundle over the algebra which carries the action of the integers.
@@
As an immediate upshot, we apply our constructions to describe the noncommutative metric geometry of quantum circle bundles while paying particular attention to the higher quantum spheres. Let us explain in more details what we mean by this statement.
@@
Another line of research, which is aligned with the present paper, seeks to construct new compact quantum metric spaces, in a functorial way, out of smaller building blocks. To our knowledge, this idea was applied for the first time in [BMR10] and has emerged again on several occasions, see e.g. [HSWZ13, KaKy22, AKK23]. Common for this approach, is an underlying decomposition of the geometry into a vertical and a horizontal direction inspired by the classical theory of fibre bundles, where the vector fields are decomposed into vertical and horizontal vector fields respecting the underlying fibre bundle structure, [BGV92, ScWa22, KavS18b]. The vertical direction of the geometry then ressembles the geometry of the fibre whereas the horizontal direction ressembles the geometry of the base space.
@@
The wording “sufficient control” is on purpose left a bit imprecise because we are currently unaware of the correct notion of a bivariant compact quantum metric space even though the paper [AKK23] presents a recent attempt at defining such a concept. Another caveat is covered by the word “spectral data” in relationship with the unbounded Kasparov product. Examples pertaining to quantum groups and crossed products by non-isometric actions show that the recipe provided by the unbounded Kasparov product does not always yield a spectral triple on the total algebra, see e.g. [KRS12, BMR10, KaKy21]. The core problem is that the horizontal Dirac operator might fail to interact with the total algebra in a way which is compatible with the bounded commutator condition of a spectral triple. Instead, it often happens that the commutator between algebra elements and the horizontal Dirac operator is displaying an exponential growth phenomenon in the vertical direction and some dampening mechanism has to be incorporated in order to tame the exponential growth.
@@ 1.1. Acknowledgements
The author gratefully acknowledge the financial support from the Independent Research Fund Denmark through grant no. 9040-00107B and 1026-00371B
@@ 3. Preliminaries on compact quantum metric spaces
It is also possible to obtain a characterization of compact quantum metric spaces in terms of finite dimensional approximations and the interested reader can consult [Kaa23, Section 3] for more details on this result.
@@ 3.2. Finitely generated projective modules
We let $L:\mathcal{X}\to[0,\infty)$ be a seminorm and $L_{0}:\mathcal{A}_{0}\to[0,\infty)$ be a slip-norm. The norm on $X$ coming from the $A_{0}$-valued inner product is denoted by $\|\cdot\|:X\to[0,\infty)$. The following theorem is a variation over [KaKy22, Theorem 2.20].
@@ 4.2. Twisted Lipschitz operators and twisted Lipschitz triples
We are now going to introduce the notion of a twisted Lipschitz operator where the commutator condition appearing in the definition of a Lipschitz operator is twisted by two particular analytic extensions of the automorphism group associated with a strongly continuous one-parameter group of unitaries. This twisting of commutators is in line with ideas appearing in [CoMo08] even though we are here insisting on a slightly more symmetric kind of twist in order to have a better relationship with the adjoint operation.
@@
To ease the notation, we put $V_{-i}:=\Gamma$ so that $V_{i}=\Gamma^{-1}$. Observe that the subspace $\textup{Dom}(\Gamma)\cap\textup{Dom}(\Gamma^{-1})\subseteq H$ is norm-dense and define the norm-dense subspace $\mathcal{H}:=(i+D)^{-1}\big{(}\textup{Dom}(\Gamma)\cap\textup{Dom}(\Gamma^{-1})\big{)}\subseteq H$. Since the resolvent $(i+D)^{-1}:H\to H$ is analytic of order $1$ and analytic of order $-1$, Proposition 4.5 entails that $\mathcal{H}\subseteq\textup{Dom}(\Gamma^{-1}D)\cap\textup{Dom}(\Gamma)$. Let now $a\in\mathcal{A}$ and apply Proposition 4.5 one more time to obtain the identities
@@ 5. First main theorem
The purpose of this section is to state and prove our first main theorem which combines the result of Li’s theorem for circle actions, Theorem 3.6, with our result on finitely generated projective modules, Theorem 3.8. We are to some extend using the setting of twisted Lipschitz triples presented in Subsection 4.2 but we do not need the full setup described there.
@@
We do not assume that $\beta_{i}$ is compatible with the involutions so that $\beta_{i}:\mathcal{A}\to B$ defined by $\beta_{-i}(a):=\beta_{i}(a^{*})^{*}$ is a unital algebra homomorphism which could be different from $\beta_{i}$.
@@
Since $P_{0}(\mathcal{A})\subseteq\mathcal{A}$ and $\mathcal{A}$ is a *-algebra we may conclude that $P_{n}(a)\in\mathcal{A}$ and hence that $P_{n}(\mathcal{A})=\mathcal{A}_{n}$. ∎
@@ 6. The vertical geometric data
As in Section 5, for each $n\in\mathbb{Z}$, we view the spectral subspace $A_{n}$ as a right Hilbert $C^{*}$-module over the fixed point algebra $A_{0}$. The relevant inner product is given by the formula $\langle a,b\rangle=a^{*}b$. We may also $A$ as a right module over $A_{0}$ and define the $A_{0}$-valued inner product given by $\langle a,b\rangle:=P_{0}(a^{*}b)$, where we recall from Subsection 3.1 that $P_{0}:A\to A$ is a faithful conditional expectation. The completion of $A$ with respect to this inner product is a right Hilbert $C^{*}$-module over $A_{0}$ which is denoted by $X$. We apply the notation $\Lambda:A\to X$ for the corresponding inclusion of $A$ into $X$. It then holds that $X$ becomes a $C^{*}$-correspondence from $A$ to $A_{0}$ with left action given by the injective unital *-homomorphism
@@ 7. The horizontal geometric data
Apply the notation $C(S^{1})$ for the unital $C^{*}$-algebra of continuous functions on the circle and let $\mathcal{O}(S^{1})\subseteq C(S^{1})$ denote the smallest unital *-subalgebra containing the inclusion $z:S^{1}\to\mathbb{C}$. Provided that the first condition in Assumption 7.1 is satisfied, we define the canonical map $\textup{can}:\mathcal{A}\otimes_{\mathcal{A}\cap A_{0}}\mathcal{A}\to\mathcal{A}\otimes\mathcal{O}(S^{1})$ by the formula $\textup{can}(x\otimes y):=\sum_{n\in\mathbb{Z}}x\cdot P_{n}(y)\otimes z^{n}$. The result formulated in the next proposition is a consequence of [AKL16, Theorem 4.3]. Notice that, in the statement, we are really referring to a quantum principal $S^{1}$-bundle for the universal differential calculus, see [Haj96, Proposition 1.6].
@@
Suppose that the conditions in Assumption 7.1 are satisfied. The $\mathbb{Z}$-graded algebra $\mathcal{A}\cong\bigoplus_{n\in\mathbb{Z}}(\mathcal{A}\cap A_{n})$ form a quantum principal $S^{1}$-bundle in the sense that the canonical map $\textup{can}:\mathcal{A}\otimes_{\mathcal{A}\cap A_{0}}\mathcal{A}\to\mathcal{A}\otimes\mathcal{O}(S^{1})$ is an isomorphism.
@@
For each $n\in\mathbb{Z}$, define the separable Hilbert space $G_{n}:=A_{n}\widehat{\otimes}_{\rho}H_{0}$ and identity $G_{n}$ with a closed subspace of $G:=X\widehat{\otimes}_{\rho}H_{0}$ via the inclusion of $A_{n}$ into $X$. Applying Lemma 6.1, we obtain a unitary isomorphism identifying $G$ with the Hilbert space direct sum ^$\widehat{\bigoplus}_{n\in\mathbb{Z}}G_{n}$.
@@
where we are suppressing the inclusion $A_{n}\widehat{\otimes}_{\rho}\mathbb{L}(H_{0})\to X\widehat{\otimes}_{\rho}\mathbb{L}(H_{0})$ for every $n\in\mathbb{Z}$. Since an infinite direct sum of essentially selfadjoint unbounded operators is again essentially selfadjoint, it follows that the associated horizontal lift
@@
is essentially selfadjoint. Indeed, upon suppressing the unitary isomorphism $G\cong\widehat{\bigoplus}_{n\in\mathbb{Z}}G_{n}$ we get the formula
@@
for all vectors $\{\xi_{n}\}_{n\in\mathbb{Z}}$ in the algebraic direct sum $\bigoplus_{n\in\mathbb{Z}}\textup{Dom}(1\otimes_{\nabla_{n}}D_{0})$ (which is now equal to $\textup{Dom}(1\otimes_{\nabla}D_{0})$).

### v2

@@ 1. Introduction
In the present paper we construct what could be called a “twisted spectral quadruple” for each of the higher quantum spheres. The key point is here that most of the quantized enveloping algebra is kept in its deformed state except for the element $K_{r}$ which is, in some sense, replaced by its classical counterpart. The element $K_{r}$ plays a special role since it is related to the circle action which describes the quantum circle bundle structure of the quantum sphere, singling out quantum projective space as invariant elements. Replacing $K_{r}$ with its classical counterpart ensures that the compact resolvent condition is satisfied for our spectral data but the price is relatively high since we are sometimes forced to consider a pair of abstract Dirac operators instead of a single one. This is precisely because of the commutator condition: One of our abstract Dirac operators interacts with the coordinate algebra via straight bounded commutators whereas the other one interacts via twisted bounded commutators. Nonetheless, both of these Dirac operators are needed in order to describe the spectral geometry of the quantum sphere since only their sum has compact resolvent, not the individual factors. Our approach is therefore related to the papers [KRS12, KaKy25, Ć24] and our work do indeed generalize parts of the results in [KaKy25].
@@
As already mentioned, we are not only focusing on the quantum spheres since our main results are formulated in the generality of quantum circle bundles. In particular, we provide general conditions which allow us to write down “twisted spectral quadruples” for quantum circle bundles. A part from the quantum spheres our methods might therefore also apply to the “quantum Grassmann spheres” which are quantum circle bundles over the quantum Grassmannians, [TaTo91, Fio99, CMOB23]. Treating this case in detail would however lead us too far astray for the time being and we therefore leave it open for future investigations. It should also be emphasized that our results to some extend cover the results of [BMR10, HSWZ13, KaKy21] regarding crossed products by the integers as one may indeed view the corresponding algebraic crossed product as a quantum circle bundle over the algebra which carries the action of the integers.
It should be clarified that our techniques are related to and inspired by the two seminal papers, [AmBa98, DaSi13], on Dirac operators on (noncommutative) circle bundles. There are nonetheless a couple of important differences which should be singled out. Contrary to the situation in [DaSi13] we do not start out with a unital spectral triple on the total algebra of a quantum circle bundle (subject to extra conditions). Our starting point is a unital spectral triple on the base algebra of a quantum circle bundle (subject to extra conditions) and the constructions carried out in this paper do not yield a unital spectral triple on the total algebra except in the special situation where a certain parameter is equal to one. In the general setting, as alluded to above, we obtain twisted spectral quadruples where the relevant twist is given in terms of an analytic extension of the circle action (at the $C^{*}$-algebraic level). For the quantum spheres the parameter just mentioned is equal to the inverse of the deformation parameter $q$ and the corresponding analytic extension of the circle action also depends on the deformation parameter $q$.
@@
As an upshot, we apply our constructions to describe the noncommutative metric geometry of quantum circle bundles while paying particular attention to the higher quantum spheres. Let us explain in more details what we mean by this statement.
@@
Another line of research, which is aligned with the present paper, seeks to construct new compact quantum metric spaces, in a functorial way, out of smaller building blocks. To our knowledge, this idea was applied for the first time in [BMR10] and has emerged again on several occasions, see e.g. [HSWZ13, KaKy25, AKK25]. Common for this approach, is an underlying decomposition of the geometry into a vertical and a horizontal direction inspired by the classical theory of fibre bundles, where the vector fields are decomposed into vertical and horizontal vector fields respecting the underlying fibre bundle structure, [BGV92, AmBa98, ScWa22, KavS18b]. The vertical direction of the geometry then resembles the geometry of the fibre whereas the horizontal direction resembles the geometry of the base space.
@@
The wording “sufficient control” is on purpose left a bit imprecise because we are currently unaware of the correct notion of a bivariant compact quantum metric space even though the paper [AKK25] presents a recent attempt at defining such a concept. Another caveat is covered by the word “spectral data” in relationship with the unbounded Kasparov product. Examples pertaining to quantum groups and crossed products by non-isometric actions show that the recipe provided by the unbounded Kasparov product does not always yield a spectral triple on the total algebra, see e.g. [KRS12, BMR10, KaKy21]. The core problem is that the horizontal Dirac operator might fail to interact with the total algebra in a way which is compatible with the bounded commutator condition of a spectral triple. Instead, it often happens that the commutator between algebra elements and the horizontal Dirac operator is displaying an exponential growth phenomenon in the vertical direction and some dampening mechanism has to be incorporated in order to tame the exponential growth.
@@ 1.1. Acknowledgements
The author gratefully acknowledge the financial support from the Independent Research Fund Denmark through grant no. 9040-00107B and 1026-00371B. This research is part of the EU Staff Exchange project 101086394 "Operator Algebras That One Can See”.
The author would also like to thank the anonymous referee for pointing out the two important papers [AmBa98, DaSi13] which we failed to reference in a first version of this paper.
1.2. Statements and Declarations
1.2.1. Conflict of interest
The author declares that there is no conflict of interest.
1.2.2. Data availability
This work uses no additional data.
@@ 3. Preliminaries on compact quantum metric spaces
It is also possible to obtain a characterization of compact quantum metric spaces in terms of finite dimensional approximations and the interested reader can consult [Kaa24, Section 3] for more details on this result.
@@ 3.2. Finitely generated projective modules
We let $L:\mathcal{X}\to[0,\infty)$ be a seminorm and $L_{0}:\mathcal{A}_{0}\to[0,\infty)$ be a slip-norm. The norm on $X$ coming from the $A_{0}$-valued inner product is denoted by $\|\cdot\|:X\to[0,\infty)$. The following theorem is a variation over [KaKy25, Theorem 2.20].
@@ 4.2. Twisted Lipschitz operators and twisted Lipschitz triples
We are now going to introduce the notion of a twisted Lipschitz operator where the commutator condition appearing in the definition of a Lipschitz operator is twisted by two particular analytic extensions of the automorphism group associated with a strongly continuous one-parameter group of unitaries. This twisting of commutators is in line with ideas appearing in [CoMo08] even though we are here insisting on a slightly more symmetric kind of twist in order to have a better relationship with the adjoint operation, see the second identity in (4.1) here below. Indeed, the symmetric kind of twist applied here is necessary for obtaining *-invariance of the associated semi-norm. This is a key condition for the theory of compact quantum metric spaces, see Definition 3.4 and the corresponding notion of a slip-norm from Definition 3.2.
@@
To ease the notation, we put $V_{-i}:=\Gamma$ so that $V_{i}=\Gamma^{-1}$. Observe that the subspace $\textup{Dom}(\Gamma)\cap\textup{Dom}(\Gamma^{-1})\subseteq H$ is norm-dense and define the norm-dense subspace $\mathcal{H}:=(i+D)^{-1}\big{(}\textup{Dom}(\Gamma)\cap\textup{Dom}(\Gamma^{-1})\big{)}\subseteq H$. Since the resolvent $(i+D)^{-1}:H\to H$ is analytic of order $1$ and analytic of order $-1$, Proposition 4.5 entails that $\mathcal{H}\subseteq\textup{Dom}(\Gamma^{-1}D)\cap\textup{Dom}(\Gamma)$. Indeed, letting $\xi\in\textup{Dom}(\Gamma)\cap\textup{Dom}(\Gamma^{-1})$ we get that $(i+D)^{-1}(\xi)\in\textup{Dom}(\Gamma)$ from Proposition 4.5 so we just need to argue that $(i+D)^{-1}(\xi)$ belongs to $\textup{Dom}(\Gamma^{-1}D)$. Clearly, $(i+D)^{-1}(\xi)$ belongs to $\textup{Dom}(D)$ and we also have that $D(i+D)^{-1}(\xi)=\xi-i(i+D)^{-1}(\xi)$ which belongs to $\textup{Dom}(\Gamma^{-1})$ by one more application of Proposition 4.5. This shows that $(i+D)^{-1}(\xi)\in\textup{Dom}(\Gamma^{-1}D)$. Let now $a\in\mathcal{A}$ and apply Proposition 4.5 to obtain the identities
@@ 5. First main theorem
The purpose of this section is to state and prove our first main theorem which combines the result of Li’s theorem for circle actions, Theorem 3.6, with our result on finitely generated projective modules, Theorem 3.8. We are to some extent using the setting of twisted Lipschitz triples presented in Subsection 4.2 but we do not need the full setup described there.
@@
We do not assume that $\beta_{i}$ is compatible with the involutions so that $\beta_{-i}:\mathcal{A}\to B$ defined by $\beta_{-i}(a):=\beta_{i}(a^{*})^{*}$ is a unital algebra homomorphism which could be different from $\beta_{i}$.
In many cases, the unital $C^{*}$-algebra $B$ will just agree with $A$ and $\beta_{i}$ (and $\beta_{-i}$) are obtained via analytic extension of the circle action $\sigma$ on $A$. This does for example happen for the quantum spheres which we are going to analyze in details later on in Section 10. However, in other cases it is necessary to work with a different choice of $B$. This situation occurs for crossed products of the form $C_{r}^{*}(\mathbb{Z},C(M))$ coming from a non-isometric diffeomorphism $\psi$ of a compact Riemannian manifold $M$, see [KaKy21, Section 4] for more details on this. But instead of diving too deep into these matters, let us continue with our exposition.
@@
Since $P_{0}(\mathcal{A})\subseteq\mathcal{A}$ and $\mathcal{A}$ is a *-algebra we may conclude that $P_{n}(a)\in\mathcal{A}$. This shows that $P_{n}(\mathcal{A})\subseteq\mathcal{A}_{n}$. The reverse inclusion follows trivially by observing that if $a\in\mathcal{A}_{n}$, then $a=P_{n}(a)$ and therefore, since $\mathcal{A}_{n}\subseteq\mathcal{A}$ we get that $a\in P_{n}(\mathcal{A})$. ∎
@@ 6. The vertical geometric data
As in Section 5, for each $n\in\mathbb{Z}$, we view the spectral subspace $A_{n}$ as a right Hilbert $C^{*}$-module over the fixed point algebra $A_{0}$. The relevant inner product is given by the formula $\langle a,b\rangle=a^{*}b$. We may also consider $A$ as a right module over $A_{0}$ and define the $A_{0}$-valued inner product given by $\langle a,b\rangle:=P_{0}(a^{*}b)$, where we recall from Subsection 3.1 that $P_{0}:A\to A$ is a faithful conditional expectation. The completion of $A$ with respect to this inner product is a right Hilbert $C^{*}$-module over $A_{0}$ which is denoted by $X$. We apply the notation $\Lambda:A\to X$ for the corresponding inclusion of $A$ into $X$. It then holds that $X$ becomes a $C^{*}$-correspondence from $A$ to $A_{0}$ with left action given by the injective unital *-homomorphism
@@ 7. The horizontal geometric data
Apply the notation $C(S^{1})$ for the unital $C^{*}$-algebra of continuous functions on the circle and let $\mathcal{O}(S^{1})\subseteq C(S^{1})$ denote the smallest unital *-subalgebra containing the inclusion $z:S^{1}\to\mathbb{C}$. Provided that the first condition in Assumption 7.1 is satisfied, we define the canonical map $\textup{can}:\mathcal{A}\otimes_{\mathcal{A}\cap A_{0}}\mathcal{A}\to\mathcal{A}\otimes\mathcal{O}(S^{1})$ by the formula $\textup{can}(x\otimes y):=\sum_{n\in\mathbb{Z}}x\cdot P_{n}(y)\otimes z^{n}$. Remark that the first condition in Assumption 7.1 does indeed imply that the sum appearing in the definition of the canonical map only has finitely many non-zero terms. The result formulated in the next proposition is a consequence of [AKL16, Theorem 4.3]. Notice that, in the statement, we are really referring to a quantum principal $S^{1}$-bundle for the universal differential calculus, see [Haj96, Proposition 1.6].
@@
Suppose that the conditions in Assumption 7.1 are satisfied. The $\mathbb{Z}$-graded algebra $\mathcal{A}\cong\bigoplus_{n\in\mathbb{Z}}(\mathcal{A}\cap A_{n})$ forms a quantum principal $S^{1}$-bundle in the sense that the canonical map $\textup{can}:\mathcal{A}\otimes_{\mathcal{A}\cap A_{0}}\mathcal{A}\to\mathcal{A}\otimes\mathcal{O}(S^{1})$ is an isomorphism.
@@
For each $n\in\mathbb{Z}$, define the separable Hilbert space $G_{n}:=A_{n}\widehat{\otimes}_{\rho}H_{0}$ and identify $G_{n}$ with a closed subspace of $G:=X\widehat{\otimes}_{\rho}H_{0}$ via the inclusion of $A_{n}$ into $X$. Applying Lemma 6.1, we obtain a unitary isomorphism identifying $G$ with the Hilbert space direct sum ^$\widehat{\bigoplus}_{n\in\mathbb{Z}}G_{n}$.
@@
where we are suppressing the inclusion $A_{n}\widehat{\otimes}_{\rho}\mathbb{L}(H_{0})\to X\widehat{\otimes}_{\rho}\mathbb{L}(H_{0})$ for every $n\in\mathbb{Z}$. Let us now argue that the associated horizontal lift
@@
is essentially selfadjoint. First of all, upon suppressing the unitary isomorphism $G\cong\widehat{\bigoplus}_{n\in\mathbb{Z}}G_{n}$ we get the formula
@@
for all vectors $\{\xi_{n}\}_{n\in\mathbb{Z}}$ in the algebraic direct sum $\bigoplus_{n\in\mathbb{Z}}\textup{Dom}(1\otimes_{\nabla_{n}}D_{0})$, which is now equal to $\textup{Dom}(1\otimes_{\nabla}D_{0})$. So, we may view $1\otimes_{\nabla}D_{0}$ as an infinite algebraic direct sum of the selfadjoint unbounded operators $1\otimes_{\nabla_{n}}D_{0}$ for $n\in\mathbb{Z}$. An application of [Sch12, Proposition 3.8] now shows that it suffices to verify that $1\otimes_{\nabla}D_{0}-i$ and $1\otimes_{\nabla}D_{0}+i$ both have dense images. But this is immediate since $1\otimes_{\nabla_{n}}D_{0}$ is selfadjoint for every $n\in\mathbb{Z}$ implying that $1\otimes_{\nabla_{n}}D_{0}-i$ and $1\otimes_{\nabla_{n}}D_{0}+i$ are in fact surjective as operators from $\mathcal{A}_{n}\otimes_{\mathcal{A}_{0}}\textup{Dom}(D_{0})$ to $A_{n}\widehat{\otimes}_{\rho}H_{0}=G_{n}$.
@@ 9. Relationship with unbounded $KK$-theory
Let us emphasize that for $\mu\neq 1$ our data does not yield a unital spectral triple over $\mathcal{A}$ (nor a twisted unital spectral triple for that matter). In fact, it appears that there is no single selfadjoint unbounded operator which adequately describes the spectral geometry of $\mathcal{A}$ in this situation. Instead, we have to work with a pair of selfadjoint unbounded operators where one of them comes from the circle action and the other one comes from the unital spectral triple over the corresponding fixed point algebra. It is therefore also unclear how to obtain a class in the $K$-homology of $A$ by taking a bounded transform since we do not know which selfadjoint unbounded operator to start out with.